OpenStudy (anonymous):

Find the inverse of the function. f(x) = 3^sqrt(x/8)- 4

OpenStudy (johnweldon1993):

Is it like \(\Large y = 3^{\sqrt{x/8}} - 4\) or \(\large y = 3^{\sqrt{x/8}-4}\) ?